OpenStudy (anonymous):

A car dealer offers you two deals on a car that costs $16,000.Please calculate the monthly payment, given these two payment options the car dealer is offering.

OpenStudy (anonymous):

please post the rest of the question

OpenStudy (anonymous):

A car dealer offers you two deals on a car that costs $16,000.Please calculate the monthly payment, given these two payment options the car dealer is offering. Payment Option 1: You can finance the car for 60 months with no interest if you make a $1,000 down payment. Payment Option 2: You can finance the car for 72 months (6 years) with 2% simple annual interest and no down payment. (Hint: To calculate simple annual interest, use the formula Interest = Principal * Rate * Time. Add the amount of interest to the price of the car.) Which monthly payment amount is lower? Please explain how you arrived at your answer and show all of your work.

OpenStudy (anonymous):

Option 1: 16,000 - 1,000 = 15,000 15,000/60 = 250 $250 per month Option 2: 16,000 (0.02) = 320/12 = 26.67/month interest 16,000/72 = 222.22/month principal 222.22 + 26.67 = $248.89 per month Option 2 is $1.11 per month lower.

OpenStudy (anonymous):

Payment Option 1: $16,000 - $1,000 = $15,000 after down payment $15,000 / 60months = $250/month Payment Option 2: $16,000 * 1.02 / 72months = $226.67/month so 2nd one is better one.

OpenStudy (anonymous):

no 1 is the correct 1